Maddie has decided to take a job as the new mascot of the MN United Loons. For her new stunt, she will stand on the lower deck and kick a ball, straight up to land in a bucket held by an assistant on the edge of the upper deck. If she kicks the ball straight up with a velocity of 24m/s and it lands in the bucket after 2.5 seconds, how high is the upper deck?
1
Expert's answer
2020-12-21T11:00:26-0500
"h=vt-0.5gt^2\\h=(24)(2.5)-0.5(9.8)(2.5)^2=29.4\ m"